Adam is a 54 year old business man who has a family history of lung cancer. Adam started smoking at 15. Years later, he got what he thought was a very bad cold that lasted for a week, started experiencing pain in his tongue, and was given ibuprofen for pain. He then noticed white sores that were getting bigger in his oral cavity near the tongue. He started feeling generally lousy and was unable to eat due to severe pain, and lost 30 lbs in a month. Adam went to the clinic, where they did a biopsy that revealed squamous cell cancer of the tongue. Adam has undergone radiation therapy (RT) and surgery to remove the cancer and a restorative surgery was scheduled. Due to RT, Adam has developed mucositis and xerostomia in his oral cavity. All his nodes that were sent to the laboratory tested negative for metastasis, and no further RT was needed. The surgeons felt confident that they had removed all the cancer. HT: 5`11” UBW: 165 lbs WT: 135 lbs Meds: Corticosteroids and antineoplastic (dexamethasone) to reduce inflammation, K depletion, Na retention; and morphine for pain.

22. What is Adam’s percent IBW? a. 78% b. 81% c. 87% d. 91% e. 103%

23. Given Adam’s weight loss, what would his MD likely diagnose him with? a. failure to thrive b. anorexia nervosa c. mildly underweight d. moderately underweight e. cancer cachexia

24. What values should be used to estimate Adam’s caloric requirements? a. 20-25 kcal/kg IBW b. 25-30 kcal/kg current body weight c. 30-35 kcal/kg current body weight d. 35-40 kcal/kg IBW e. 40-45 kcal/kg IBW
Answer #1

22. B) 81%

ibw = 50+[11*2.3]= 75.2

wt in kg = 61.235

%wt in ibw = 61.235/75.2*100 = 81%

23) c mild underweight

24) b. 25-30 kcal/kg current body weight